# Probing displaced (dark)photons from low reheating freeze-in at the LHC

### v1

@@ 3.1 Boltzmann equations
where $H/\overline{H}\equiv 1+\frac{T}{3}\frac{d\log g_{*s}}{dT}$, the Hubble rate is given by $H=\sqrt{\frac{\pi^{2}g_{*}}{90}}\frac{T^{2}}{M_{P}}$, where $g_{*}(T)$ is the effective number of relativistic degrees of freedom at temperature $T$, and $M_{P}=2.4\times 10^{18}$ GeV is the reduced Planck mass. The $C$ factors in the right side of each BE correspond to integrated collision terms, with $C_{A\leftrightarrow B}\equiv C_{A\rightarrow B}-C_{B\rightarrow A}$, with Belanger:2018ccd
@@
where in eq. (15) we have used again detailed balance in order to express both collision terms as a function of $\expectationvalue{\sigma v}_{\phi\phi\rightarrow f\bar{f}}$. We are thus allowed to use the algebraic non-relativistic expansion of $\expectationvalue{\sigma v}_{\phi\phi\rightarrow ff}$ given by Cline:2013gha
@@
where $\Gamma_{h}\approx 4$ MeV, and $X_{f}$ is a factor equal to 1 for leptons and $\approx 3$ for quarks, including one-loop QCD corrections (see appendix A of Cline:2013gha ). Similarly to the case in eq. (14) involving the dark photon portal for freeze-in, expression (17) is useful for computational calculations.
@@
Before presenting the numerical results obtained from solving the coupled BEs (7) and (8), we must emphasize two things. Firstly, the results in the subsequent sections are only reliable within the parameter space we focus on; for instance, for much higher $T_{\rm RH}$, additional collision terms in the coupled Boltzmann equations—such as $W^{+}W^{-}$ annihilation—should be included (see e.g. Arias:2025nub ). Secondly, the non-relativistic expansion used to compute $\expectationvalue{\sigma v}_{\gamma\textquoteright\phi\rightarrow f\bar{f}}$ becomes invalid for $T_{\rm RH}\gg m_{\gamma^{\prime}}$ as $\gamma\textquoteright$ ’ behaves relativistically in the cosmic frame. Notice that in the region of parameter space where the dark photon portal significantly contributes to the collision terms of the coupled Boltzmann equations (e.g., for relatively small values of $\lambda_{HS}$), the non-relativistic approximation can differ by up to a factor of two in the calculation of the relic abundance, compared to the standard result obtained by thermally averaging $\sigma v$ Gondolo:1990dk .
@@ 3.2 Results
Figure 3: Yield evolution of $\phi$ (gray lines) and $\gamma^{\prime}$ (black lines) starting from $T_{\rm RH}=2~{}\mathrm{GeV}$, considering $g_{D}=10^{-11}~{}\mathrm{GeV}^{-1}$ (left panel) and $10^{-9}~{}\mathrm{GeV}^{-1}$ (right panel). In both plots, we fix $m_{\gamma^{\prime}}=1~{}\mathrm{GeV}$ and $m_{\phi}=30~{}\mathrm{GeV}$. The red and orange solid lines correspond to the equilibrium yields, while the red dotted line in the right panel represents the combination $Y_{\gamma^{\prime}}\,Y_{\phi,\mathrm{eq}}/Y_{\gamma^{\prime},\mathrm{eq}}$ for the case $\lambda_{HS}=0.1$. The red dotted line illustrates the quasi-static equilibrium Hall:2009bx , where $\phi$ participates via the process $\phi\leftrightarrow\gamma^{\prime}\gamma$ before full chemical decoupling.
@@
  • Thermalized $\phi$: For larger portal coupling $\lambda_{HS}$ (solid gray lines), $\phi$ can thermalize with the SM plasma ($Y_{\phi}\approx Y_{\phi,e}$). After thermal decoupling, it decays contributing to the relic abundance of $\gamma^{\prime}$ (in this case, $\gamma^{\prime}$ behaves as a super-WIMP Feng:2003xh ; Feng:2003uy ; Asaka:2005cn ; Asaka:2006fs ; Molinaro:2014lfa ; Faber:2019mti ). For small $g_{D}$ (left panel), this delayed production significantly enhances $Y_{\gamma^{\prime}}$. For larger $g_{D}$ (right panel), the decay of $\phi$ is too fast to impact $\gamma^{\prime}$ significantly. [^6] In this regime, $\phi$ decouples from equilibrium before fully depleting, entering a so-called quasi-static equilibrium. This behavior is typical of mediators that decay faster than they annihilate, and may eventually freeze-out.
[…]
@@
[…]
@@ 3.3 Cosmological constraints of the parameter space
The cosmological history of the present model is strongly influenced by the lifetime of $\phi$ and its abundance prior to decay. Particularly dangerous are decays occurring around the epoch of light element formation, BBN. On the one hand, electromagnetic decays can modify the baryon-to-photon ratio and disintegrate light nuclei. Additionally, if the decay occurs after neutrino decoupling, it can lower the neutrino temperature. On the other hand, decays into relativistic dark photons can impact BBN through contributions to the effective number of neutrino species. Moreover, we must ensure that the dark photon cools sufficiently before matter-radiation equality, so as not to interfere with the formation of large-scale structures in the early Universe. Since these constraints depend only on the $g_{D}$ coupling, they have already been studied in Arias:2025nub ; nevertheless, we briefly review them here for completeness.
@@ 3.3.1 Dark matter warmness
In order not to interfere with the formation of large structures in the Universe, the DM velocity is restricted to satisfy Viel:2005qj ; Viel:2007mv ; Baur:2015jsy ; Chatrchyan:2020pzh $v_{\rm DM}\lesssim 10^{-3}c,\,\mbox{at}\,\,a=a_{\rm EQ}$, with $a_{\rm EQ}$ the scale factor at matter-radiation equality. Dark photons from the decay of $\phi$ gain substantial momentum, given by
@@ 3.3.2 BBN constraints
A detailed treatment of the above scenarios would require solving the full set of BBN equations, including the evolution of light nuclei. Since such an analysis goes beyond the scope of this work, we instead rely on the results of Ref. Yeh:2024ors , where the authors study the cosmological impact of a general non-relativistic species $X$ decaying near the BBN epoch, accounting for both visible and invisible decay modes.
@@
In scenarios where the decay products are electromagnetic, Ref. Yeh:2024ors finds a mild preference for a non-zero energy injection, driven in part by a slight preference in CMB observations for $N_{\nu}<3$. The preferred region corresponds to $\xi(\tau_{X}/1\,\text{sec})^{1/2}\sim 0.015$ for lifetimes between $1$ and $100$ seconds. On the other hand, in models with invisible (dark sector) decays, the strongest constraints arise from $N_{\rm eff}$, with BBN data setting the leading limits for short lifetimes and CMB data dominating at larger lifetimes.
Our scenario interpolates between these limiting cases. Decays into photons can lower $N_{\rm eff}$ and affect the baryon asymmetry, while invisible decays can enhance the radiation density. To ensure consistency with observational constraints, we conservatively adopt the bound $\xi\lesssim 10^{-5}$, assume an initial temperature of $T_{\rm ini}=10$ MeV before BBN, and restrict the lifetime of $\phi$ to below $10^{4}$ seconds, which is totally safe for our purposes. This ensures that the decays occur before the Universe cools below $T=10$ keV, thereby avoiding severe constraints from the photodisintegration of light elements Cyburt:2002uv ; Hufnagel:2017dgo ; Hufnagel:2018bjp ; Kawasaki:2020qxm ; Depta:2020zbh .
@@ 4.1.1 Bound from non-pointing photons search
The analysis motivating this work uses the full Run 2 data set of proton-proton collisions delivered by the LHC at a center-of-mass energy of $\sqrt{s}=13$ TeV between 2015 and 2018 and recorded by the ATLAS detector, corresponding to an integrated luminosity of $139~{}{\rm fb}^{-1}$. The experiment searches for delayed and non-pointing photons, which would originate from the displaced decay of a neutral long-lived particle (LLP) ATLAS:2022vhr . The measurement is triggered by a prompt electron or muon coming from associated production with the Higgs. The results are interpreted in a scenario in which the LLPs are produced in pairs in exotic decays of the $125$ GeV Higgs boson, with each LLP subsequently decaying into a photon and a particle that escapes direct detection, giving rise to missing transverse momentum. No significant excess is observed above the expectation due to background processes, so upper limits are set on the branching ratio of the exotic decay of the Higgs boson. In addition, a model-independent limit is also provided for the production cross-section of photons with large values of displacement and time delay.
@@
The displacement and time delay of the photon are both reconstructed through the electromagnetic shower of the photon within the ATLAS Electromagnetic Calorimeter (ECal). The time delay $t_{\gamma}$ is defined as the relative arrival time to the ECal, with respect to what is expected from a prompt photon. Furthermore, the non-pointing parameter $|\Delta z_{\gamma}|$ is defined as the separation between the extrapolated origin of the photon and the position of the primary vertex of the event, projected onto the beamline in the $z-r$ plane, as depicted in figure 7. Both variables are determined in the first two layers of the ECal, with the $|\Delta z_{\gamma}|$ measurement also being possible due to the varying lateral segmentation of the detector, which allows reconstruction of the flight direction of the photon. Then, for prompt decays, both $|\Delta z_{\gamma}|$ and $t_{\gamma}$ are naively expected to be zero, meaning that these are useful handles for identifying neutral long-lived particles decaying into photons.
Let us now briefly describe the event selection procedure in the experimental analysis. [^7] A more detailed description of the recast can be found in Delgado:2022fea ; Duarte:2023tdw . The technical details of the current refinements on the simulation of our signal events, their analysis and statistical interpretation are described in appendix B. The photons and charged leptons had to satisfy selection cuts, as described in Duarte:2023tdw . In the search, in addition to the lepton trigger that requires a prompt electron or muon with $p_{T}>27$ GeV, the selected events must have at least one photon candidate, being divided into two orthogonal final states, with the $1\gamma$ final state requiring exactly one photon (single-photon channel) and the $\geq 2\gamma$ final state requiring two or more photons (multiphoton channel). Regardless of this classification, the analysis uses the information $t_{\gamma}$ and $|\Delta z_{\gamma}|$ from the photon with the largest $p_{\gamma T}$ in the barrel, on which a cut $E_{\rm cell}>10$ GeV is applied ($E_{\rm cell}$ is defined as the maximum energy deposit of the shower in the middle layer of the ECal). Events are also required to have a minimum value of missing transverse energy ${\rm MET}>50$ GeV to account for the escaping particles, in our case, the dark photons and neutrinos from leptonic decays of the associated system.
In order to contribute to any of the aforementioned channels, the analyzed photon must satisfy $|\Delta z_{\gamma}|\,>300$ mm. Moreover, for the single-photon channel, the search requires $121.5<t_{\gamma}<$ 12 ns, while the multi-photon channel needs $121<t_{\gamma}<$ 12 ns. The number of events measured in these bins, as well as the background, is given in table II of ATLAS:2022vhr .
@@
[…]
@@ 4.1.2 Bound from Higgs invisible decays
Higgs invisible decays refer to situations where the Higgs boson decays either to stable neutral particles or to very long-lived neutral particles which decay outside the detector. Such decays are reflected on $B_{\rm inv}$, the branching ratio of the Higgs into the invisible final state. The strongest limit in the literature to this branching ratio comes from a combination of direct searches, performed by ATLAS ATLAS:2023tkt , reading $B_{\rm inv}<0.107$.
@@
Thus, we have $\gamma_{\rm rel}\,\beta_{\rm rel}$ between $6.2-0.29$ for $m_{\phi}$ between $10-60$ GeV, respectively. These can be interpreted as the minimum values for $\gamma_{\rm rel}\,\beta_{\rm rel}$ when the Higgs is boosted.
In the following, we shall compare our bounds from non-pointing photon searches with a conservative limit from Higgs invisible decays. For this specific constraint, we will indeed assume that the Higgs decays at rest and that the ATLAS ECal is spherical, with $L_{\rm det}=1.97$ m ATLAS:2008xda . This implies:

### v2

@@ 3.1 Boltzmann equations
where $H/\overline{H}\equiv 1+\frac{T}{3}\frac{d\log g_{*s}}{dT}$, the Hubble rate is given by $H=\sqrt{\frac{\pi^{2}g_{*}}{90}}\frac{T^{2}}{M_{P}}$, where $g_{*}(T)$ is the effective number of relativistic degrees of freedom at temperature $T$, and $M_{P}=2.4\times 10^{18}$ GeV is the reduced Planck mass. The $C$ factors in the right side of each BE correspond to integrated collision terms, with $C_{A\leftrightarrow B}\equiv C_{A\rightarrow B}-C_{B\rightarrow A}$, with Bélanger et al. (2018)
@@
where in eq. (15) we have used again detailed balance in order to express both collision terms as a function of $\expectationvalue{\sigma v}_{\phi\phi\rightarrow f\bar{f}}$. We are thus allowed to use the algebraic non-relativistic expansion of $\expectationvalue{\sigma v}_{\phi\phi\rightarrow ff}$ given by Cline et al. (2013)
@@
where $\Gamma_{h}\approx 4$ MeV, and $X_{f}$ is a factor equal to 1 for leptons and $\approx 3$ for quarks, including one-loop QCD corrections (see appendix A of Cline et al. (2013)). Similarly to the case in eq. (14) involving the dark photon portal for freeze-in, expression (17) is useful for computational calculations.
@@
We have implemented the set of BEs considering the approximations presented above, and cross-checked the relevant cross sections and decay widths with micrOMEGAs. Before presenting the numerical results, we emphasize the following. Firstly, the results in the subsequent sections are only reliable within the parameter space we focus on; for instance, for much higher $T_{\rm RH}$, additional collision terms in the coupled Boltzmann equations—such as $W^{+}W^{-}$ or $hh$ annihilation—should be included (see e.g. Arias et al. (2025)). Secondly, the non-relativistic expansion used to compute $\expectationvalue{\sigma v}_{\gamma^{\prime}\phi\rightarrow f\bar{f}}$ becomes invalid for $T_{\rm RH}\gg m_{\gamma^{\prime}}$ as $\gamma^{\prime}$ behaves relativistically in the cosmic frame. Notice that in the region of parameter space where the dark photon portal significantly contributes to the collision terms of the coupled Boltzmann equations (e.g., for relatively small values of $\lambda_{HS}$), the non-relativistic approximation can differ by up to a factor of two in the calculation of the relic abundance, compared to the standard result obtained by thermally averaging $\sigma v$ Gondolo and Gelmini (1991).
@@ 3.2 Results
Figure 3: Yield evolution of $\phi$ (gray lines) and $\gamma^{\prime}$ (black lines) starting from $T_{\rm RH}=2~\mathrm{GeV}$. In both plots, we fix $m_{\gamma^{\prime}}=1~\mathrm{GeV}$ and $m_{\phi}=30~\mathrm{GeV}$. The red and orange solid lines correspond to the equilibrium yields, while the red dotted line in the right panel represents the combination $Y_{\gamma^{\prime}}\,Y_{\phi,\mathrm{eq}}/Y_{\gamma^{\prime},\mathrm{eq}}$ for the case $\lambda_{HS}=0.1$. The red dotted line illustrates the quasi-static equilibrium Hall et al. (2010), where $\phi$ participates via the process $\phi\leftrightarrow\gamma^{\prime}\gamma$ before full chemical decoupling.
@@
  • Thermalized $\phi$: For larger portal coupling $\lambda_{HS}$ (solid gray lines), $\phi$ can thermalize with the SM plasma ($Y_{\phi}\approx Y_{\phi,e}$). After thermal decoupling, it decays contributing to the relic abundance of $\gamma^{\prime}$ (in this case, $\gamma^{\prime}$ behaves as a super-WIMP Feng et al. (2003a, b); Asaka et al. (2006, 2007); Molinaro et al. (2014); Faber et al. (2020)). For small $g_{D}$ (left panel), this delayed production significantly enhances $Y_{\gamma^{\prime}}$. For larger $g_{D}$ (right panel), the decay of $\phi$ is too fast to impact $\gamma^{\prime}$ significantly. [^7] In this regime, $\phi$ decouples from equilibrium before fully depleting, entering a so-called quasi-static equilibrium. This behavior is typical of mediators that decay faster than they annihilate, and may eventually freeze-out.
To illustrate the differences between the thermal and non-thermal $\phi$ scenarios in a more general way, in figure 4 we show the result of a random scan over a selected region of the parameter space for $m_{\gamma^{\prime}}=1$ GeV, retaining only the points that yield the correct relic abundance $\Omega_{\gamma^{\prime}}h^{2}=0.12\pm 0.02$. The scan is performed using a linear scale for $m_{\phi}\in[10,\,62.5]$ GeV and $T_{\rm RH}\in[1,\,5]$ GeV, and a logarithmic scale for $g_{D}\in[10^{-11},\,10^{-9}]$ GeV-1 and $\lambda_{HS}\in[10^{-3},\,10^{-1}]$. In the figure, the black points correspond to the scenarios in which $\phi$ thermalizes and undergoes freeze-out, while for the orange points $\phi$ never enters in equilibrium with the SM.
@@
  • As $g_{D}$ decreases and $T_{\rm RH}$ increases, the freeze-in production of $\phi$ via the Higgs portal interaction becomes relevant, and dominates the production of $\gamma^{\prime}$ indirectly. The vertical line in the $g_{D}$$T_{\rm RH}$ plane is straightforward to interpret: since the dark masses and Higgs portal coupling are fixed, requiring the correct relic abundance determines $T_{\rm RH}$, independently of $g_{D}$ (which does not affect the yield in this regime).
@@
In the $\lambda_{HS}$$T_{\rm RH}$ plane (right panel of figure 5), we see that when $g_{D}$ is fixed, reproducing the correct relic abundance at larger $\lambda_{HS}$ requires lowering $T_{\rm RH}$. The super-WIMP production appears in the upper part of the plot ($\lambda_{HS}\sim 0.1$) as a plateau in $T_{\rm RH}$ for the dashed lines (corresponding to smaller $g_{D}$). In this regime, the $\phi$ yield is set by freeze-out, thus independent of $T_{\rm RH}$. However, as $T_{\rm RH}$ increases, $\phi$ decays efficiently while in equilibrium, interrupting the freeze-out and altering the relic abundance.
@@ 3.3 Cosmological constraints of the parameter space
The cosmological history of the present model is strongly influenced by the lifetime of $\phi$ and its abundance prior to decay. Particularly dangerous are decays occurring around the epoch of light element formation, BBN. On the one hand, electromagnetic decays can modify the baryon-to-photon ratio and disintegrate light nuclei. Additionally, if the decay occurs after neutrino decoupling, it can lower the neutrino temperature. On the other hand, decays into relativistic dark photons can impact BBN through contributions to the effective number of neutrino species. Moreover, we must ensure that the dark photon cools sufficiently before matter-radiation equality, so as not to interfere with the formation of large-scale structures in the early Universe. Since these constraints depend only on the $g_{D}$ coupling, they have already been studied in Arias et al. (2025); nevertheless, we briefly review them here for completeness.
@@ 3.3.1 Dark matter warmness
In order not to interfere with the formation of large structures in the Universe, the DM velocity is restricted to satisfy Viel et al. (2005, 2008); Baur et al. (2016); Chatrchyan and Jaeckel (2021) $v_{\rm DM}\lesssim 10^{-3}c,\,\mbox{at}\,\,a=a_{\rm EQ}$, with $a_{\rm EQ}$ the scale factor at matter-radiation equality. Dark photons from the decay of $\phi$ gain substantial momentum, given by
@@ 3.3.2 BBN constraints
A detailed treatment of the above scenarios would require solving the full set of BBN equations, including the evolution of light nuclei. Since such an analysis goes beyond the scope of this work, we instead rely on the results of Ref. Yeh et al. (2024), where the authors study the cosmological impact of a general non-relativistic species $X$ decaying near the BBN epoch, accounting for both visible and invisible decay modes.
@@
In scenarios where the decay products are electromagnetic, Ref. Yeh et al. (2024) finds a mild preference for a non-zero energy injection, driven in part by a slight preference in CMB observations for $N_{\nu}<3$. The preferred region corresponds to $\xi(\tau_{X}/1\,\text{sec})^{1/2}\sim 0.015$ for lifetimes between $1$ and $100$ seconds. On the other hand, in models with invisible (dark sector) decays, the strongest constraints arise from $N_{\rm eff}$, with BBN data setting the leading limits for short lifetimes and CMB data dominating at larger lifetimes.
Our scenario interpolates between these limiting cases. Decays into photons can lower $N_{\rm eff}$ and affect the baryon asymmetry, while invisible decays can enhance the radiation density. To ensure consistency with observational constraints, we conservatively adopt the bound $\xi\lesssim 10^{-5}$, assume an initial temperature of $T_{\rm ini}=10$ MeV before BBN, and restrict the lifetime of $\phi$ to below $10^{4}$ seconds, which is totally safe for our purposes. This ensures that the decays occur before the Universe cools below $T=10$ keV, thereby avoiding severe constraints from the photodisintegration of light elements Cyburt et al. (2003); Hufnagel et al. (2018a, b); Kawasaki et al. (2020); Depta et al. (2021).
@@ 4.1.1 Bound from non-pointing photons search
The analysis motivating this work uses the full Run 2 data set of proton-proton collisions delivered by the LHC at a center-of-mass energy of $\sqrt{s}=13$ TeV between 2015 and 2018 and recorded by the ATLAS detector, corresponding to an integrated luminosity of $139~{\rm fb}^{-1}$. The experiment searches for delayed and non-pointing photons, which would originate from the displaced decay of a neutral long-lived particle (LLP) Aad and others (2023b). The measurement is triggered by a prompt electron or muon coming from associated production with the Higgs. [^8] We note that Higgs production via gluon fusion could also be taken into account for this search by using a single-jet trigger instead, requiring $p^{j}_{T}>100-200$ GeV Aaboud and others (2017). We have estimated that, at leading order, the corresponding cross-section is of an order of magnitude similar to those for production with an associated lepton. The results are interpreted in a scenario in which the LLPs are produced in pairs in exotic decays of the $125$ GeV Higgs boson, with each LLP subsequently decaying into a photon and a particle that escapes direct detection, giving rise to missing transverse momentum. No significant excess is observed above the expectation due to background processes, so upper limits are set on the branching ratio of the exotic decay of the Higgs boson. In addition, a model-independent limit is also provided for the production cross-section of photons with large values of displacement and time delay.
@@
The displacement and time delay of the photon are both reconstructed through the electromagnetic shower of the photon within the ATLAS Electromagnetic Calorimeter (ECal). The time delay $t_{\gamma}$ is defined as the relative arrival time to the ECal, with respect to what is expected from a prompt photon. Furthermore, the non-pointing parameter $|\Delta z_{\gamma}|$ is defined as the separation between the extrapolated origin of the photon and the position of the primary vertex of the event, projected onto the beamline in the zrz-r plane, as depicted in figure 7. Both variables are determined in the first two layers of the ECal, with the $|\Delta z_{\gamma}|$ measurement also being possible due to the varying lateral segmentation of the detector, which allows reconstruction of the flight direction of the photon. Then, for prompt decays, both $|\Delta z_{\gamma}|$ and $t_{\gamma}$ are naively expected to be zero, meaning that these are useful handles for identifying neutral long-lived particles decaying into photons.
Let us now briefly describe the event selection procedure in the experimental analysis. [^9] A more detailed description of the recast can be found in Delgado et al. (2022); Duarte et al. (2024). The technical details of the current refinements on the simulation of our signal events, their analysis and statistical interpretation are described in appendix B. The photons and charged leptons had to satisfy selection cuts, as described in Duarte et al. (2024). In the search, in addition to the lepton trigger that requires a prompt electron or muon with $p_{T}>27$ GeV, the selected events must have at least one photon candidate, being divided into two orthogonal final states, with the $1\gamma$ final state requiring exactly one photon (single-photon channel) and the $\geq 2\gamma$ final state requiring two or more photons (multiphoton channel). Regardless of this classification, the analysis uses the information $t_{\gamma}$ and $|\Delta z_{\gamma}|$ from the photon with the largest $p_{\gamma T}$ in the barrel, on which a cut $E_{\rm cell}>10$ GeV is applied ($E_{\rm cell}$ is defined as the maximum energy deposit of the shower in the middle layer of the ECal). Events are also required to have a minimum value of missing transverse energy ${\rm MET}>50$ GeV to account for the escaping particles, in our case, the dark photons and neutrinos from leptonic decays of the associated system.
In order to contribute to any of the aforementioned channels, the analyzed photon must satisfy $|\Delta z_{\gamma}|\,>300$ mm. Moreover, for the single-photon channel, the search requires $1.5<t_{\gamma}<12$ ns, while the multi-photon channel needs $1<t_{\gamma}<12$ ns. The number of events measured in these bins, as well as the background, is given in table II of Aad and others (2023b).
@@ 11 $2+$ 11 $2+$
Table 1: Cutflow showing efficiency (in $\%$) for the displaced photon search, taking $m_{\phi}=60$ GeV, $m_{\gamma^{\prime}}=1$ GeV. Here, we combine ZhZh, $W^{\pm}h$ and $t\bar{t}h$ production processes. We show results for $g_{D}=5.7\times 10^{-10}$ and $1.6\times 10^{-10}$ GeV-1, which respectively imply decay lengths $c\,\tau_{\phi}=0.69$ and $4.4$ m. We distinguish between single ($1$) and multi- ($2+$) photon channels.
@@ 4.1.2 Bound from Higgs invisible decays
Higgs invisible decays refer to situations where the Higgs boson decays either to stable neutral particles or to very long-lived neutral particles which decay outside the detector. Such decays are reflected on $B_{\rm inv}$, the branching ratio of the Higgs into the invisible final state. The strongest limit in the literature to this branching ratio comes from a combination of direct searches, performed by ATLAS Aad and others (2023a), reading $B_{\rm inv}<0.107$.
@@
Thus, we have $\gamma_{\rm rel}\,\beta_{\rm rel}$ between 6.20.296.2-0.29 for $m_{\phi}$ between 106010-60 GeV, respectively. These can be interpreted as the minimum values for $\gamma_{\rm rel}\,\beta_{\rm rel}$ when the Higgs is boosted.
In the following, we shall compare our bounds from non-pointing photon searches with a conservative limit from Higgs invisible decays. For this specific constraint, we will indeed assume that the Higgs decays at rest and that the ATLAS ECal is spherical, with $L_{\rm det}=1.97$ m Aad and others (2008). This implies: